# Diffusion Reward: Learning Rewards via Conditional Video Diffusion

### v2

@@ 4.2 Conditional Entropy as Rewards
Estimation of conditional entropy. To formalize this idea, we seek to estimate the negative conditional entropy given historical frames $\bm{z}_{c}$, which in principle captures the conditional generation diversity:
@@
One primary challenge is the computation of the entropy in Eq. (1), which is intractable as we have no explicit form of the conditional distribution [34]. Therefore, we instead attempt to estimate the variational bound of such entropy. Specifically, we first present the variational bound of conditional log-likelihood as follows:
@@
where $\bm{z}_{k}^{0}$ is the denoised prediction of current observation $\bm{z}_{k}$. This bound could be estimated via noise predictor $\bm{\epsilon}_{\theta}$ [15, 33], or with the closed-form distribution [32, 20] (e.g., discrete multivariate distribution). We use the latter one as our choice of estimation because of its better compatibility with VQ-Diffusion.
@@
We visualize the aggregated reward in Figure 4 and present curves of each task in the Appendix. The results show that conditional entropy can successfully capture the varied generative diversity on different videos, echoed with the aforementioned insight of Diffusion Reward. Notably, here we use the standardized entropy reward $\bar{r}^{\mathrm{ce}}$ to mitigate the burden of hyperparameter tuning, as we observe that the scale of conditional entropy varies significantly across different tasks and domains, partially attributed to the varied objects and environment dynamics. Concretely, the conditional entropy is standardized by the empirical mean and standard deviation of the expert videos:
@@
Exploration reward. As the reward $\bar{r}^{\mathrm{ce}}$ incentivizes the agent to mimic the behavioral patterns of the expert, the exploration may still be prohibitively challenging in complex tasks with high-dimensional input. To alleviate this issue, we incorporate RND [3] as the exploration reward, termed as $r^{\mathrm{rnd}}$.
Diffusion Reward. To this end, we combine our proposed diffusion-based entropy reward with the exploration reward and the raw sparse reward $r^{\mathrm{spar}}$ from the environment,
$r^{\mathrm{diff}}=(1-\alpha)\cdot\bar{r}^{\mathrm{ce}}+\alpha\cdot r^{\mathrm{rnd}}+r^{\mathrm{spar}},$ (5)
where $\alpha$ is the reward coefficient that steers the weight of two rewards. The integration of $r^{\mathrm{spar}}$ is crucial, as the complete absence of environmental supervision may hinder progress in tackling complex tasks.
@@
Figure 5: Task visualization. We evaluate our method on 10 challenging visual RL tasks from MetaWorld and Adroit, with $64\times 64$-dimensional RGB images and sparse rewards. Tasks are chosen to cover a wide range of manipulation skills.
4.3 Training Details
We first provide the expert videos from various tasks as the whole dataset for pertaining reward models, which can be generated by scripted policies or other means. Then, we first use VQ-GAN [10] to train the encoder, associated with $8\times 8$ size of codebook across all domains, with an additional perceptual loss [43] calculated by a discriminator to increase perceptual quality. We subsequently use VQ-Diffusion [12] for training the conditional video diffusion model, where the number of condition frames is set as 2 for all tasks. For reward inference, the reward coefficient is set as $0.95$ for all tasks except 00 for the Pen task. Moreover, during the diffusion process, we also use a DDIM-like sampling strategy [33] to accelerate the diffusion process with denoise steps set as $10$ and repeated with $1$ times for all tasks, which shows fair performance and retains high inference speed.
Figure 6: Main results. Success rate for our method and baselines on 7 gripper manipulation tasks from MetaWorld and 3 dexterous manipulation tasks from Adroit with image observations. Our method achieves prominent performance on all tasks, and significantly outperforms baselines on complex door and hammer tasks. Results are means of 5 runs with standard errors (shaded area).
5 Experiments
In this section, we first introduce the overall experimental setups in Section 5.1. We then present the significant performance improvement of our method against competitive baselines in Section 5.2, investigate the generalization capability and effectivenss on real robot videos of our learned reward in Section 5.3 and Section 5.4, respectively. We also conduct comprehensive ablations to study the effect of each component and implementation details in Section 5.5.
5.1 Experimental Setup
@@
Simulation envrionments. We intend to demonstrate the effectiveness of Diffusion Reward on 10 complex visual robotic manipulation tasks, including 7 gripper manipulation tasks from MetaWorld [40] and 3 dexterous hand manipulation tasks from Adroit [28], as visualized in Figure 10. We choose these two simulation environments because of their task diversity and complexity. Each task is associated with $64\times 64$-dimensional RGB images, ±$\pm 4$ pixel shift augmentation [39], and 0/1 environmental sparse reward. For pertaining reward models, we collect 20 expert videos for each Metaword task via the scripted policy provided by the official repository, and 50 for Adroit via the policies trained with performant RL method [36]. For downstream RL training, the interaction budget of the Adroit task is set as $3$ million to ensure convergence of using our rewards, and is respectively set for the MetaWorld task due to the varied task complexity.
Baselines. We compare our method against the followings:
@@
  • Raw Sparse Reward that uses the environmental sparse reward. This comparison tests the benefit of adding our pre-trained reward.
@@
  • Random Network Distillation (RND, [3]) that encourages exploration with a novelty-seeking reward. This comparison tests the benefit of rewarding the agent with expert-like behaviors.
@@
  • Adversarial Motion Priors (AMP, [27]) that learns a discriminator to discern agent behaviors and expert behaviors based on current observations. This comparison tests the benefit of encoding temporal information in learned reward and utilization of novelty-seeking reward.
@@
  • Video Prediction Rewards (VIPER, [9]) and its standardized variant (VIPER-std.) that use VideoGPT [38] as video prediction model and predicted log-likelihood of agent observation as reward. This comparison tests the benefit of utilizing the generative capability of video diffusion models and the conditional entropy as a more explorative reward.
@@
For a fair comparison, all methods use DrQv2 [39] as the RL backbone and maintain all settings except reward-pretraining (if exist) identical.
Figure 7: Success rate curves on 5 unseen MetaWorld tasks. Diffusion Reward could generalize to unseen tasks directly and produce reasonable rewards, largely surpassing other baselines. Results are means of 4 runs with standard errors (shaded area).
@@ 5.2 Main Results
Comparison with non-pretraining methods. We present the learning curves of success rates for each method over two simulation domains in Figure 1 and 6. The results show that solely using sparse task rewards enables progress in relatively straightforward tasks such as Reach and Dial Turn. However, it encounters significant challenges in more complex ones, exemplified by the Door and Hammer within the dexterous hand manipulation domain. The incorporation of pure novelty-seeking rewards (RND) unsurprisingly enhances the RL agent’s exploration, particularly evident in addressing moderately complex tasks like Coffee Push and Assembly. Nevertheless, the performance on dexterous hand manipulation tasks is still unsatisfactory due to the lack of expert-instructed exploration at prohibitively large configuration space. Conversely, AMP explicitly employs expert-guided rewards to incentivize the exploration of expert-like behaviors. While it generally outperforms RND in simpler tasks, its efficacy diminishes in more complex ones such as Lever Pull and Hammer.
Comparison with reward pretraining method. The above observations suggest that the combination of expert-instructed rewards and novelty-seeking rewards is likely to perform favorably. Despite the incorporation of such a combination in the VIPER reward, its empirical performance unexpectedly falls short of both RND and AMP. We posit this is attributed to the significantly varied scales of expert-instructed rewards (i.e., log-likelihood), which may eliminate the effect of novelty-seeking rewards. VIPER-std alleviates this issue and shows better performance than the non-standardized version, especially on Coffee Push and Pen, while still underperforming our method. These results are in line with two limitations outlined in Section 4.2 and further verified in Figure 2, indicating VIPER’s struggle in capturing complex video distributions within intricate tasks. In sharp contrast to VIPER, our proposed method not only utilizes the modeling capabilities of diffusion models, but also uses conditional entropy as a reward function to accelerate exploration. The results showcase remarkable performance improvements of 38% and 35% over the best-performing baselines with the same training steps across MetaWorld and Adroit, respectively.
@@ 5.3 Zero-Shot Reward Generalization
Reward visualization. Video diffusion models have exhibited powerful abilities for modeling videos and, notably, for generating samples beyond their training data, as exemplified in text-to-image video generation [16, 11]. Such an advance motivates our exploration into the potential of Diffusion Reward to generalize to previously unseen tasks. To investigate this, we start by visualizing the learned returns of trajectories with varying qualities derived from 15 unseen gripper manipulation tasks in MetaWorld (see Figure 4). Interestingly, the distinctions between trajectories of different qualities are less noticeable in comparison to those observed in tasks seen during training. Nevertheless, our pre-trained reward model still exhibits a consistent trend where expert-like behaviors receive relatively higher learned returns, owing to the generalization prowess of the video diffusion model.
Figure 8: Reward curve of real robot videos. Our method assigns higher rewards to expert videos than to random videos.
RL performance. Subsequently, we directly apply our pre-trained rewards to 5 tasks involving diverse objects without additional tuning. The outcomes, as shown in Figure 7, affirm that our reward effectively guides RL exploration and largely outperforms other baseline methods across all tasks. Notably, our approach proves helpful in constraining the exploration space of RND due to its retained ability to discriminate between expert-like and expert-unlike behaviors. In contrast, VIPER struggles to generalize effectively on most tasks, partially attributed to the combined limitations of the adopted video models and log-likelihood rewards. These results not only verify the efficacy of our method, but also point towards the potential of employing larger diffusion models and integrating other modalities (e.g., text embedding for task specification) to enhance the generalization capabilities of our approach further.
Figure 9: Ablations. Success rate curves for ablated versions of Diffusion Reward, aggregated over Door and Hammer from Adroit. (a) We test the combinations of generative models and rewards to show the benefit of estimating conditional entropy with the diffusion model. (b) We ablate the choice of the number of denoising steps. (c) We demonstrate that the inherent randomness of Diffusion Reward from the reverse process helps RL exploration. (d) We ablate the choice of reward coefficient. (e) We test the effect of the number of conditional frames. Results are means of 3 seeds with standard error (shaded area). Red is our default.
@@ 5.4 Real Robot Evaluation
We consider the real robot task that aims to pick up a bowl on the table. To train our reward model, we collect 20 real robot videos (10 expert and 10 random) with an Allegro hand, a Franka arm, and a RealSense. Visualization results in Figure 8 show that Diffusion Reward can appropriately assign expert videos relatively higher rewards and, in contrast, random videos lower rewards, indicating the potential of our method for real-world robot manipulation tasks. We present more curves of multiple trials in the Appendix.

### v3

@@ 4.1 Expert Video Modeling via Diffusion Model
Figure 4: Time efficiency on Door task with a NVIDIA A40 GPU.
Latent diffusion process.
Specifically, we first train an encoder unsupervisedly from expert videos to compress the high-dimensional observations. Here we use the VQ-GAN method [12] to represent the image $\bm{x}$ with a vector-quantized latent code $\bm{z}=Q(E(\bm{x}))$, where $E$ is the encoder and $Q$ is the element-wise quantizer. The whole video is then represented by a sequence of latent variables $\tau=\{\bm{z}_{0},\bm{z}_{1},...,\bm{z}_{K-1}\}$, where we overwrite the definition of $\tau$ without ambiguity. Subsequently, the forward process applies noise $\epsilon$ in the latent space at each time step $t\in 0,...,T$ to the data distribution $\bm{z}_{k}$, resulting in a noisy sample $\bm{z}_{k}^{t}$, where $\bm{z}_{k}^{t}=\sqrt{\bar{\alpha}_{t}}\bm{z}_{k}^{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$, and $\bar{\alpha}$ is the accumulation of the noise schedule over past timesteps. To fit data distribution, we learn a parameterized variant of noise predictor $\bm{\epsilon}_{\theta}(\bm{z}_{k}^{t})$ that aims to predict the noise $\bm{\epsilon}$ during the forward process. Then the parameterized reverse process $p_{\theta}(\bm{z}_{k}^{t-1}|\bm{z}_{k}^{t})$ can be approximated and performed by iterative denoising an initial distribution.
Historical frames as condition.
To utilize the temporal information from expert videos with the power of video diffusion, we further condition the reverse process with historical frames, i.e., $p_{\theta}(\bm{z}_{k}^{t-1}|\bm{z}_{k}^{t},\bm{z}_{c})$, where $\bm{z}_{c}$ is the concatenation of $\ell$ historical frames $[\bm{z}_{k-\ell},...,\bm{z}_{k-1}]$ to ensure higher computation efficiency while maintaining enough historical information. This can also be viewed as matching the distribution of expert and agent trajectories [10]. Subsequently, one can perform the reverse process from a randomly sampled noise to generate the latent code of future frames and decode the code for video prediction. In this work, we use VQ-Diffusion [13] as our choice of video diffusion model due to its good performance and compatibility with vector-quantized latent code, but our framework can in principle adopt any off-the-shelf video diffusion models. Specifically, we first tokenize each latent code $\bm{z}$ indexed by its indices that specify the respective entry in the learned codebook of VQ-GAN, and take as a condition embedding the concatenated tokens. The embedding is then fed to the decoder that contains multiple transformer blocks and softmax layers with cross attention, providing context information for video diffusion.
4.2 Conditional Entropy as Rewards
While previous studies have explored the use of log-likelihood as rewards with video prediction models, exemplified by works such as VIPER [10], this approach encounters two primary challenges. Firstly, it has difficulty in accurately modeling the distribution of complex expert videos featuring intricate dynamics, as shown in Table 2. Secondly, the moderate video modeling ability leads to undesired rewards. This issue is evident in Figure 3, where a significant drop in learned rewards between in-distribution expert videos and out-of-distribution ones, though both sets of videos demonstrate optimal behaviors.
Key insight behind Diffusion Reward.
We address these challenges by harnessing the great generative capability of video diffusion models. Our observations indicate increased generation diversity with unseen historical observations (rand.) and reduced diversity with seen ones (expert), as shown in Table 2. This gives rise to the key insight of our proposed Diffusion Reward: diffusion conditioned on expert-like trajectories exhibits lower diversity where the agent ought to be rewarded more, and the opposite holds on unexpert-unlike ones. Such diversity discrimination incentivizes RL agents to chase expert-like behaviors.
Estimation of conditional entropy.
To formalize this idea, we seek to estimate the negative conditional entropy given historical frames $\bm{z}_{c}$, which in principle captures the conditional generation diversity:
@@ 4.2 Conditional Entropy as Rewards
One primary challenge is the computation of the entropy in Eq. (1), which is intractable as we have no explicit form of the conditional distribution [44]. Therefore, we instead estimate the variational bound of the entropy. Specifically, we first present the variational bound of conditional log-likelihood as follows:
@@
where $\bm{z}_{k}^{0}$ is the denoised prediction of current observation $\bm{z}_{k}$. This bound can be estimated via noise predictor $\bm{\epsilon}_{\theta}$ [16, 43], or with the closed-form distribution [42, 20] (e.g., discrete multivariate distribution). We use the latter one as our choice of estimation because of its better compatibility with VQ-Diffusion.
@@
We visualize the aggregated entropy-based reward in Figure 3 and present curves of each task in the Appendix. The results show that conditional entropy can successfully capture the varied generative diversity on different videos, echoed with the aforementioned insight of Diffusion Reward. Notably, here we use the standardized entropy reward $\bar{r}^{\mathrm{ce}}$ to mitigate the burden of hyperparameter tuning, as we observe that the scale of conditional entropy varies significantly across different tasks and domains, partially attributed to the varied objects and environment dynamics. Concretely, the conditional entropy is standardized by the empirical mean and standard deviation of the expert videos:
@@
Exploration reward.
As the reward $\bar{r}^{\mathrm{ce}}$ incentivizes the agent to mimic the behavioral patterns of the expert, the exploration may still be prohibitively challenging in complex tasks with high-dimensional input [10, 1]. To alleviate this issue, we incorporate RND [4] as the exploration reward, termed as $r^{\mathrm{rnd}}$.
Diffusion Reward.
Lastly, we combine our proposed diffusion-based entropy reward with the exploration reward and the raw sparse task reward $r^{\mathrm{spar}}$ from the environment, yielding the following composite reward function:
$r^{\mathrm{diff}}=(1-\alpha)\cdot\bar{r}^{\mathrm{ce}}+\alpha\cdot r^{\mathrm{rnd}}+\beta\cdot r^{\mathrm{spar}},$ (5)
where $\alpha$ and $\beta$ are the reward coefficients that steer the effect of each reward. The inclusion of $r^{\mathrm{spar}}$ is crucial, as the complete absence of environmental supervision can impede progress in tackling complex tasks. Furthermore, we opt for a single coefficient $\alpha$ to balance $\bar{r}^{\mathrm{ce}}$ and $r^{\mathrm{rnd}}$, thereby alleviating the burden of hyperparameter tuning. Meanwhile, the coefficient of task reward $\beta$ is set as 1 for all tasks. We present negligible empirical distinctions between this formulation and the one featuring an additional coefficient, and effect of $\beta$ in the Appendix.
4.3 Training Details
We first provide the expert videos from various tasks as the whole dataset for pertaining reward models, which can be generated by scripted policies or other means. Then, we first use VQ-GAN [12] to train the encoder, associated with $8\times 8$ size of codebook across all domains, with an additional perceptual loss [56] calculated by a discriminator to increase perceptual quality. We subsequently use VQ-Diffusion [13] for training the conditional video diffusion model, where the number of historical frames $\ell$ is set as 2 for all tasks. For reward inference, the reward coefficient is set as $0.95$ for all tasks except 00 for the Pen task. We decrease $\alpha$ for the Pen task because a high exploration reward can misguide the robot hand to drop the pen, while a lower one encourages in-hand manipulation. Moreover, during the diffusion process, we also use a DDIM-like sampling strategy [43] to accelerate the diffusion process with denoise steps set as $10$ and repetition of diffusion $M$ as $1$ for all tasks, which shows fair performance and retains high inference speed in Figure 4. More details on hyperparameters, network architectures, and downstream RL can be found in the Appendix.
5 Experiments
5.1 Experimental Setup
@@
Figure 5: Task visualization. We evaluate methods on 10 challenging visual RL tasks from MetaWorld and Adroit with visual input and sparse rewards. Tasks are chosen to cover a wide range of manipulation skills.
Figure 6: Reward curve of real robot videos. Our method assigns higher rewards to expert videos than to random ones.
Simulation environments.
@@
We intend to demonstrate the effectiveness of Diffusion Reward on 10 complex visual robotic manipulation tasks, including 7 gripper manipulation tasks from MetaWorld [52] and 3 dexterous hand manipulation tasks from Adroit [36], as visualized in Figure 12. We choose these two simulation environments because of their task diversity and complexity. Each task is associated with $64\times 64$-dimensional RGB images, ±$\pm 4$ pixel shift augmentation [51], and 0/1 sparse task reward. For pertaining reward models, we collect 20 expert videos for each Metaword task (5039 frames in total) via the scripted policy provided by the official repository, and 50 for Adroit (3346 frames in total) via the policies trained with performant RL method [47], while our method has emerging low demand for the data scale (see the Appendix). For downstream RL training, the interaction budget of the Adroit task is set as $3$ million to ensure convergence of using our rewards, and is respectively set for the MetaWorld task due to the noticeably varied task complexity.
Baselines.
We compare our method against the followings:
@@
  • Raw Sparse Reward that uses the sparse task reward. This comparison tests the benefit of adding learned rewards from videos.
@@
  • Random Network Distillation (RND, [4]) that encourages exploration with a novelty-seeking reward only. This comparison tests the benefit of rewarding the agent with our pretrained reward.
@@
  • Adversarial Motion Priors (AMP, [33]) that online learns a discriminator to discern agent behaviors and expert behaviors based on current observations. This comparison tests the benefit of encoding temporal information in learned reward and utilization of novelty-seeking reward.
@@
  • Video Prediction Rewards (VIPER, [10]) and its standardized variant (VIPER-std) that use VideoGPT [50] as video prediction model and predicted log-likelihood of agent observation as reward. This comparison tests the benefit of utilizing the generative capability of video diffusion models and the conditional entropy as a more explorative reward.
@@
For a fair comparison, all methods are augmented with task rewards, which generally improve their performance. Besides, we use DrQv2 [51] as the RL backbone and maintain all settings except reward-pretraining (if exist) identical.
Figure 7: Main results. Success rate for our method and baselines on 7 gripper manipulation tasks from MetaWorld and 3 dexterous manipulation tasks from Adroit with image observations. Our method achieves prominent performance on all tasks, and significantly outperforms baselines on complex door and hammer tasks. Results are means of 5 runs with the standard error (shaded area).
@@ 5.2 Main Results
Comparison with non-pretraining methods.
We present the learning curves of success rates for each method over two simulation domains in Figure 1 and 7. The results show that solely using sparse task rewards enables progress in relatively straightforward tasks such as Reach and Dial Turn. However, it encounters significant challenges in more complex ones, exemplified by the Door and Hammer within the dexterous hand manipulation domain. The incorporation of pure novelty-seeking rewards (RND) unsurprisingly enhances the RL agent’s exploration, particularly evident in addressing moderately complex tasks like Coffee Push and Assembly. Nevertheless, the performance on dexterous hand manipulation tasks is still unsatisfactory due to the lack of expert-instructed exploration at prohibitively large configuration space. Conversely, AMP explicitly employs expert-guided rewards to incentivize the exploration of expert-like behaviors. While it generally outperforms RND in simpler tasks, its efficacy diminishes in more complex ones such as Lever Pull and Hammer.
Comparison with reward pretraining method.
The above observations suggest that the combination of expert-instructed rewards and novelty-seeking rewards is likely to perform favorably. Despite the incorporation of such a combination in the VIPER reward, its empirical performance unexpectedly falls short of both RND and AMP. We posit this is attributed to the significantly varied scales of expert-instructed rewards (i.e., log-likelihood), which may eliminate the effect of novelty-seeking rewards. VIPER-std alleviates this issue and shows better performance than the non-standardized version, especially on Coffee Push and Pen, while still underperforming our method. These results are in line with two limitations outlined in Section 4.2 and further verified in Figure 3, indicating VIPER’s struggle in capturing complex video distributions within intricate tasks. In sharp contrast to VIPER, our proposed method not only utilizes the modeling capabilities of diffusion models, but also uses conditional entropy as a reward function to accelerate exploration. The results showcase remarkable performance improvements of 38% and 35% over the best-performing baselines with the same training steps across MetaWorld and Adroit, respectively.
Figure 8: Success rate curves on 5 unseen MetaWorld tasks. Diffusion Reward could generalize to unseen tasks directly and produce reasonable rewards, largely surpassing other baselines. Results are means of 4 runs with the standard errors (shaded area).
@@ 5.3 Zero-shot Reward Generalization
Reward analysis.
The advances of video diffusion models in generating samples beyond their training data, as exemplified in text-to-image video generation [15, 11], motivates our exploration into the potential of Diffusion Reward to generalize to previously unseen tasks. To investigate this, we first visuale the learned returns of trajectories with varying qualities derived from 15 diverse unseen tasks from MetaWorld (see Figure 3). Expectedly, the distinctions between trajectories of varied qualities are less noticeable than those observed in tasks seen during pretraining. Nonetheless, our pretrained reward model still exhibits a consistent trend where expert-like behaviors receive relatively high learned returns, owing to the generalization prowess of the video diffusion model.
RL performance.
Subsequently, we directly apply our pretrained rewards to 5 tasks (more in the Appendix) involving diverse objects without additional tuning. The outcomes, as shown in Figure 8, affirm that our reward effectively guides RL exploration and largely outperforms other baseline methods across all tasks. Notably, our approach proves helpful in constraining the exploration space of RND due to its retained ability to discriminate between expert-like and expert-unlike behaviors. Meanwhile, VIPER generalizes worse than ours on most tasks, especially on complex ones like opening objects, partially attributed to the combined limitations of the adopted video models and log-likelihood rewards. Here we exclude comparison with AMP because it learns reward online, which requires unavailable expert data on the target task. These results not only verify the efficacy of our method, but also point towards the potential of employing larger diffusion models and integrating other modalities (e.g., text-based task specification) to enhance the generalization capabilities of our approach further.
Figure 9: Generalization in real-world task. Policy trained with our method can successfully press the button at an unseen position.
Figure 10: Success rate of real-world offline RL. Our method significantly outperforms baselines on the Button Press task given different offline data.
@@ 5.4 Real Robot Evaluation
Reward analysis.
We consider the real robot task that aims to pick up a bowl on the table. To train and test our reward model, we collect 20 real robot videos (10 expert and 10 random) with an Allegro hand, a Franka arm, and a RealSense, operated by humans. Visualization results in Figure 6 show that Diffusion Reward can appropriately assign expert videos relatively higher rewards and, in contrast, random videos lower rewards, indicating the potential of our method for real-world robot manipulation tasks. We present more curves of multiple trials and reward analysis related to baselines in the Appendix.
Offline RL performance.
We conduct offline RL evaluations in the Button Press task using a parallel gripper as the end-effector. Expert videos are used to train each reward model for subsequent offline RL training (see the Appendix for more details). Our results show that Diffusion Reward produces a more generalizable policy for pressing the button at an unseen position, as illustrated in Figure 10. Moreover, Figure 10 quantitatively demonstrates that Diffusion Reward significantly outperforms VIPER in terms of real-world applicability.